adult tickets for the school concert cost $5 and student tickets cost $3.50. the school sold $2,162.50 worth of tickets. if 120 more students than adults attended the cocert, how many students attended the concert?
Answer by mananth(16949)

let the number of adults be x
students will be x+120
5x+3.5(x+120)=2162.5
5x+3.5x+420=2162.5
8.5x=2162.50-420
8.5x=1742.50
x=1742.50/8.5
=205 adults
120+205 =325 srudents attended